convert the following equation written in cartesian coordinate in to an equation of cylindrical coordinates x^3-2x^2-6z=4-2y^2

1
Expert's answer
2021-11-08T08:51:55-0500

"x^3 -2x^2+2y^2 -6z-4=0"

Put x = r cos (A)

and, y = r sin (A)

and z =z, we get

"(r cos A)^3 - 2 (r cos A)^2 + 2 (r sin A)^2 - 6z -4=0"

"r^3 cos^3A - 2r^2 cos^2A + 2r^2 sin^2A - 6z-4=0"

"r^3cos^3A - 2r^2 (cos^2A-sin^2A) - 6z-4=0"

"r^3cos^3A -2 r^2cos2A - 6z-4=0"


Final Cylindrical equation:

"r^3cos^3A - 2r^2cos2A - 6z-4=0"